You have 20 coins some dimes and some nickels and you need to make $1.40 but you need more dimes than nickels
Answer by macston(5194)

.
D=dimes; N=nickels
D+N=20
D=20-N
$0.10D+$0.05N=$1.40 Substitute for D
$0.10(20-N)+$0.05N=$1.40
$2.00-$0.10N+$0.05N=$1.40 Subtract $2.00 from each side.
-$0.05N=-$0.60 Divide each side by $0.05
N=12 ANSWER 1: There are 12 nickels
D=20-N=20-12=8 ANSWER 2: There are 8 dimes.
CHECK:
$0.10D+$0.05N=$1.40
$0.10(8)+$0.05(12)=$1.40
$0.80+$0.60=$1.40
$1.40=$1.40